een the two buildings is D, and the rouilding is 2.0 m below the jumping-of pint. Find the mamum value for D. at a speed of 5.9 m/s, hop to land on the roof of an adjacent building. Air resistance is GO TUTORIAL Question Attempts: Unlimited SAVE FOR LATER MapleNet
Answer #1

using kinematics

y = vyt - 1/2gt2

as there is no vertical component of speed, vy = 0

y = - 1/2gt2

so,

t = sqrt (2y / g)

t = 0.638876 sec

Therefore

D = vx * t

D = 5.9 * 0.638876

D = 3.769 m